there are 3 liquids a,b,c of equal masses,they are at temprature10,15,25centigrade.when liquid a and b are mixed together common temprature is 12centigrade when b and c are mixed then common temprature is21centigrade.what will be the common temprature of a and c.


 

Grade:

1 Answers

Anurag Kishore
13 years ago

Hi

Use principle of calorimetry

For A and B

m sa (12 - 10) = m sb (15 - 12)

=> sa / sb = 3/2   ---------------------------------(1)

 

For B and C

m sb (21 - 15) = m sc (25 - 21)

=> sb / sc = 4/6 = 2/3 --------------------------------(2)

From (1) and (2)

sa / sc = 3/2*2/3 = 1  -----------(3)

 

Now when A and C mixed let final temp = T

m sa (T - 10) = m sc (25 - T)                          [from(3)]

=> T =35/2 = 17.5 0C

 

 

 

 

Thanks

Anurag Kishore